OpenStudy (darkbluechocobo):

Match the use of fuel in miles per gallon to the car or the truck. 1. truck driving in the city 2. truck driving on the highway 3. car driving in the city 4. car driving on the highway a. 28 mi/gal b. 26 mi/gal c. 24 mi/gal d. 35 mi/gal

OpenStudy (darkbluechocobo):

@Tyler1992 Alright could you help me with this one. This makes no sense at all

OpenStudy (anonymous):

well a truck is going to get less miles to the gallon than a car because its bigger. So the 24 and 26 are for the truck and the other two are for the car.

OpenStudy (anonymous):

Now if you're driving on the highway its most likely you're not speeding up or slowing down, but driving at constant speed so your miles per gallon will be better on the highway than in the city.

OpenStudy (anonymous):

So out of the two we said were for the truck, the bigger of the two will be for the highway and the other one will be for the city. Same for the car

OpenStudy (darkbluechocobo):

O sorry I didn't see you replied

OpenStudy (darkbluechocobo):

@tyler1992 Would you say that 26 would be for a highway and a 24 would be a city?

OpenStudy (anonymous):

for the truck, yes

OpenStudy (darkbluechocobo):

So now for the second part how would you find this How many hours of highway driving are required for each to empty its tank? Assume an average speed of 60 miles per hour. The truck will have an empty tank after 7.93 h. The car will have an empty tank after 6.07 h. The truck will have an empty tank after 6.8 h. The car will have an empty tank after 6.07 h. The truck will have an empty tank after 6.8 h. The car will have an empty tank after 8.17 h. The truck will have an empty tank after 7.93 h. The car will have an empty tank after 8.17 h.

OpenStudy (darkbluechocobo):

I was thinking that the truck would have more gas in it because a truck is ment to go long distances without stopping for gas but thats just me

OpenStudy (anonymous):

It doesn't tell you how big their tank is?

OpenStudy (darkbluechocobo):

I am assuming you have to take the data that was given in the first question and put it into this question

OpenStudy (darkbluechocobo):

so basically take the highway driving the the mpg would be

OpenStudy (anonymous):

But you have to know how big their tank is so you know how many miles they can go. They could have a 15 gallon tank or a 29 gallon tank.

OpenStudy (anonymous):

because i believe you use the formula \[speed = \frac{distance}{time}\]

OpenStudy (darkbluechocobo):

so then it would be D if I am correct

OpenStudy (anonymous):

How'd you come to that conclusion? Can you show me?

OpenStudy (darkbluechocobo):

For the second part, you just divide the total number of miles by 60 to get the length in hours. For the truck, it's 476/60 = 7.9333 For the car, it's 490/60 = 8.1666

OpenStudy (anonymous):

where did you get the total miles from?

OpenStudy (darkbluechocobo):

With the 1st part of the question To figure out the miles per gallon, you put the number of miles driven on a full tank over the gallons in a tank. Truck: 408/17 City = 24 mpg in the city, 476/17 Hwy = 28 mpg hwy Car: 364/14 city = 26 mpg in the city, 490/14 hwy = 35 mpg

OpenStudy (anonymous):

right, i just never saw the total miles driven on a full tank. Im asuming it was given to you in the problem?

OpenStudy (darkbluechocobo):

Well Yes i just saw that

OpenStudy (anonymous):

Oh ok, i just wasn't sure if i was missing something obvious lol :P but yeah thats what you do so you're correct!

OpenStudy (darkbluechocobo):

But thank you for the help :3

OpenStudy (anonymous):

No problem!